y= - sqrt(x-1)+1
graph and find the domain and range

Answer by stanbon(48564)

y= - sqrt(x-1)+1
graph and find the domain and range
--------------
The x-1 must be >=0
So, x must be >= 1
Domain: All Real numbers >=1
---------
The sqrt(x+1) must be >=0
So, -sqrt(x+1) must be <=0
And, -sqrt(x+1)+1 must be <=1
Range: All Real numbers <=1
========
Cheers,
Stan H.